OpenStudy (teamus1021):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4 f(n -1) - 3 f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -2, 8, -26, -80, -2, 8, -26, 80, -2, 12, -44, 180, -2, -12, -44, -180,

OpenStudy (eliesaab):

It is either a or b. Why?

OpenStudy (teamus1021):

i need help

OpenStudy (eliesaab):

did you understand why the answer is a or b?

OpenStudy (teamus1021):

@newwar

OpenStudy (teamus1021):

no

OpenStudy (eliesaab):

They gave you the first two terms of the sequence. What are they?

OpenStudy (teamus1021):

negative and positive

OpenStudy (eliesaab):

what numbers are they?

OpenStudy (teamus1021):

-80 and 80

OpenStudy (teamus1021):

-2, 8, -26, -80,

OpenStudy (eliesaab):

f(0)=? f(1)=?

OpenStudy (teamus1021):

-2, 8, -26, 80

OpenStudy (teamus1021):

how do i figure out if its negative and positive

OpenStudy (eliesaab):

what make you decide that?

OpenStudy (teamus1021):

im confused

OpenStudy (eliesaab):

-2, 8, -26, 80

OpenStudy (eliesaab):

is the answer

OpenStudy (teamus1021):

thank u

OpenStudy (teamus1021):

i have one more can u help

OpenStudy (teamus1021):

The first four terms of a sequence are shown below. 8, 5, 2, -1 Which of the following functions best defines this sequence? (4 points) f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (teamus1021):

@eliesaab

OpenStudy (teamus1021):

@jigglypuff314

OpenStudy (eliesaab):

You should always ask a new problem in a new post. Look 8, 5, 2, -1 5=8-3 2=5-3 -1=2-3 What does this tell you?

OpenStudy (eliesaab):

Each term is equal the term before it -3, starting from the second one.

OpenStudy (eliesaab):

Which of the choices describe this phenomenon